Your firm successfully issued new debt last​ year, but the debt carries covenants.​ Specifically, you can only pay dividends out of earnings made after the debt issue and you must maintain a minimum quick​ (acid-test) ratio

left parenthesis Current Assets minus Inventory right parenthesis divided by Current Liabilities(Current Assets−Inventory)/Current Liabilities

of 1.1. Your net income this year was $70.5 million. Your cash is $9.6 ​million, your receivables are $7.8 million, and your inventory is $5.4 million. You have current liabilities of $19.3 million. What is the maximum dividend you could pay​ (in cash and in​ stock) this year and still comply with your​ covenants?

The maximum dividend would be

​$nothing

million.  ​(Round to one decimal​ place.)
Answer #1

Quick Ratio = 1.1

=> (Current Assets - Inventory) / Current Liabilities = 1.1

=> Current Assets = 1.1 x Current Liabilities + Inventory = 1.1 x 19.3 + 5.4 = $26.63 million

=> Cash + Receivables + Inventory - Current Liabilities + Net Income - Dividends = 26.63

=> Max Dividends = 9.6 + 7.8 + 5.4 - 19.3 + 70.5 - 26.63 = $37.37 million